OpenStudy (anonymous):

Solve the equation: 5x(x + 3) = 0.

OpenStudy (saifoo.khan):

x = 0 x = -3

OpenStudy (anonymous):

5x(x + 3) = 0. 5x=0 x+3=0 x=-3

OpenStudy (anonymous):

5x(x + 3) = 0. 5x=0 x+3=0 x=-3 x=0